OpenStudy (anonymous):

1. Solve the equation 10.5 = 5z a. -1/2 b. 15.5 c. 2.1 d. 5 2. Solve the equation -9 = 5/17n a. -153/5 b. -5/153 c. -17/45 d. -45/17

OpenStudy (anonymous):

10.5=5z z=10.5/2=2.1 -9=5/17n n=-5/(17*9)=-5/153

OpenStudy (anonymous):

u wanna get z by itself so u divide 5 on both side so u get z= 2.1 for the second 1 u divide 5/17 on both sides so u get 5/153

OpenStudy (anonymous):

okay thanks for explaining :)